Camp Rainbow offers overnight summer camp programs for children ages 10–14 every summer during June and July. Each camp session is one week and can accommodate up to 200 children. The camp is not coed, so boys attend during the odd-numbered weeks and girls attend during the even-numbered weeks. While at the camp, participants make crafts, participate in various sports, help care for the camp’s resident animals, have cookouts and hayrides, and help assemble toys for local underprivileged children.

The camp provides all food as well as materials for all craft classes and the toys to be assembled. One cabin can accommodate up to 10 children, and one camp counselor is assigned to each cabin. Three camp managers are on-site regardless of the number of campers enrolled.

Following is the cost information for Camp Rainbow’s operations last summer:

Week Number of Campers Cost to Run Camp
1 178 $13,000
2 126 9,010
3 118 9,980
4 138 11,500
5 90 8,370
6 108 8,880
7 140 11,800
8 190 14,360


Required:
4. Using the high-low method, calculate Camp Rainbow’s total fixed operating costs and variable operating cost per child. (Round your intermediate calculations and final answer to 2 decimal places.)

Fixed Cost
Variable Cost per Child





5. Using the high-low method results, calculate the camp’s expected operating cost if 180 children attend a session. (Round your intermediate calculations and final answer to 2 decimal places.)

Expected Cost:
Answer #1
Camp Rainbow
4.Using the high-low method, calculate Camp Rainbow’s total fixed operating
costs and variable operating cost per child
$14,360-$8,370/ 190-90 = $59.90 = Variable cost/camper
5. Using the high-low method results, calculate the camp’s expected operating
cost if 180 children attend a session.
High- $14,360 – (190 x $59.90) = $14,360 - $11,381 = $2,979 = Fxed cost
Low- $8,370 – (90 x $59.90) = $8,370 - $5,391 = $2,979
Y= a + bx
Y= $2,979 +$59.90(x)= $2,979 +$59.90(180)=$13,761 (estimate)